An object displaced from 5m height reaches ground at a distance of 10m.what is its final velocity?
1
Expert's answer
2019-08-20T10:24:08-0400

For horizontal projectile motion the distance


"d=v_0\sqrt{\frac{2h}{g}}"

So, the initial velocity

 

"v_0=d\sqrt{\frac{g}{2h}}=10\sqrt{\frac{10}{2\times 5}}=10\:\rm{m/s}"